# AI4EOSC: a Federated Cloud Platform for Artificial Intelligence in Scientific Research

### 17 Mar 2026 v2

@@ Abstract
In this paper, we describe the architecture and implementation of a federated compute platform dedicated to support Artificial Intelligence and Machine Learning in scientific workloads. Putting the effort into reproducible deployments, it delivers consistent, transparent access to a federation of physically distributed e-Infrastructures, with a particular focus on the European Open Science Cloud. Through a comprehensive service catalogue, the platform is able to offer an integrated user experience covering the full Machine Learning lifecycle, including model development (with dedicated interactive development environments), training (with GPU resources, annotation tools, experiment tracking, and federated learning support) and deployment (covering a wide range of deployment options all along the Cloud Continuum). The platform also provides tools for traceability and reproducibility of AI models, integrates with different Artificial Intelligence model providers, datasets and storage resources, allowing users to interact with the broader Machine Learning ecosystem. Finally, it is easily customizable to lower the adoption barrier by external communities.
@@
Cloud Platform , Artificial Intelligence , Machine Learning , European Open Science Cloud , MLOps
@@ 1 Introduction
Artificial Intelligence (AI), Machine Learning (ML), and Deep Learning (DL) are transforming scientific and industrial research. The impact of these techniques over the last years has revolutionized the field in a wide range of applications—e.g., by making it possible to tackle new scientific endeavors [21, 48, 31], by automatizing manual and tedious tasks [51] or by speeding up some time-consuming and and effort-intensive tasks [85], just to cite some examples. These techniques are a reality, and they form nowadays part of the basic set of tools for a wide set of scientists and researchers [42].
However, this data-driven research revolution demands more than just access to computational and storage resources. Scientists willing to effectively exploit these state-of-the-art techniques require advanced platforms and tools that can lower adoption barriers (e.g., access to accelerators), easing and simplifying data exploration and analysis, model development, training, sharing and deployment into production environments; in an open, transparent and collaborative way. All these aspects go beyond simple model development and require addressing complex engineering challenges that span the entire AI/ML development lifecycle.
In a partial response to these challenges, Machine Learning Operations (MLOps) has emerged as a set of practices and tools designed to ease the management of the entire AI/ML lifecycle [9], analogous to how DevOps bridges the gap between software development and IT operations. MLOps focuses on model deployment automation, reproducibility, and scalability, facilitating the seamless transition from data science experimentation to the reliable operation of ML components in production. While MLOps has been successfully adopted by industry, its adoption in the scientific and research context is not that prominent. Many existing MLOps platforms lack important open-science-centric features that scientists require nowadays, such as seamless integration with research infrastructures and data repositories, or direct support for FAIR principles [91].
@@
This gap is particularly critical in the context of the European Open Science Cloud (EOSC) ecosystem. The EOSC is an European Commission initiative aimed at creating the “web of FAIR data and services”, as a transformative process towards the promotion of open science practices among researchers, also delivering a comprehensive environment for the sharing, processing, and analysis of research data [24]. In the context of AI/ML applications for scientific research, realizing this vision requires not only infrastructure but also platforms that can operationalize open science principles across the entire AI/ML lifecycle.
In this context, we present the AI4EOSC platform, a federated system that provides a comprehensive cloud-based toolbox for the development, training, sharing and deployment of AI/ML solutions, with a strong focus on traceability, reproducibility, and open science principles. AI4EOSC enables data processing from the cloud-to-edge continuum, being able to mobilize different resources for different computational needs. Moreover, being developed in the context of the EOSC ecosystem, the platform delivers seamless technical and semantical interoperability, ensuring and strengthening open science practices as a collaborative, transparent, and accessible research model.
The AI4EOSC platform is not merely focused on providing scientists with the tools and functionalities needed to develop, train or serve models, but also to frame the development of these models with the open science practices —and hence, the FAIR principles [91]— by design, ensuring fine-grained model provenance metadata generation our automated metadata interoperability across different formats. To this aim, we have implemented these functionalities in an automated and transparent way for the users, so that these principles can be granted by design, to the possible extent, without their intervention.
AI4EOSC operationalizes and adopts MLOps in a research context, extending its principles to support the unique needs of open science. The platform provides scientists with the robust, reproducible, and scalable workflows they need, ensuring compliance with FAIR principles and comprehensive integration with the EOSC and European Research Infrastructures ecosystem.
@@
The remainder of this manuscript is structured as follows: first, Section 2 introduces the background and related work. In Section 3 and Section 4, we discuss the general architecture of the AI4EOSC platform and we introduce further considerations that underpin the global design and architecture respectively. In Section 5 we explain how different scientists with multidisciplinary profiles can leverage the platform to carry out different tasks. Finally, Section 6 draws out the main conclusions and the future work proposed.
@@ 2 Related work and motivation
While there are many private and proprietary platforms to deploy AI/ML workloads for different kind of users, few of them have put an effort on addressing the open science and reproducibility requirements that are key to scientific workloads. There is a clear framework fragmentation, where the existing solutions address a specific part of the AI/ML lifecycle, often being isolated. Moreover, we argue that there is a clear lack of integrated and open-science focused solutions that span the whole AI/ML lifecycle.
2.1 Existing solutions: a fragmented ecosystem
2.1.1 Model development, training and reproducibility
The foundational phase in AI/ML research is the model development and training. This process is normally supported by a wide range of frameworks and tools like Scikit-learn, JAX, PyTorch or TensorFlow [63], providing the needed functionality for model creation, with a strong focus on performance and scalability. These tools normally require access to accelerators (e.g., GPU, TPU, NPU) in order to be able to efficiently build an AI/ML model.
Experiment tracking platforms, including open source —e.g., MLflow [93] or LabML [45]— and proprietary ones —like Weights and Biases [25], Valohai [88] or Comet [23]— allow developers to monitor and compare training runs, but their focus remains on industry workflows, with limited support for open science practices (due to the lack of structured metadata in some cases) or the lack of integration with research infrastructures. Similarly, reproducibility platforms like Code Ocean  [^1] https://codeocean.com/ or Binder [49] allow the creation of reproducible environments, but these are typically isolated from the broader AI/ML lifecycle, lacking connections to model sharing or deployment.
@@ 2.1.1 Model development, training and reproducibility
There are some AI/ML platforms that do cover the whole lifecycle process end-to-end. However, these are closed source (e.g., H2O.ai [41]) and restrict model deployment or the data storage to their specific cloud solutions (like Google Vertex AI [4], Amazon Sage Maker [6] or Azure ML Studio [58]), thus defeating the open principles on which science is built. Others are partially open-source but do not offer all features in their community free versions (Polyaxon [72]).
Beyond these commercial solutions, the DEEP-Hybrid-DataCloud project proposed a distributed architecture to provide ML developers with a set of cloud services that allowed them to ease the model creation, training, validation, and model serving [56]. The DEEP platform focused on easing the access to accelerators (GPUs) in pan-European e-Infrastructures, but it lacked technical or semantical interoperability with other systems, especially due to the recent developments and advancements in the EOSC federation and the European Research Infrastructures.
2.1.2 Model sharing and model repositories
A number of AI model repositories exist, including those domain-specific that are devoted to scientific communities, like the BioImage Model Zoo [66] for microscopy images or Kipoi [7] for genomics. Also, community-specific model repositories like BioModels [39, 57] are being augmented, in order to include AI/ML assets along with mathematical models of biological and biomedical systems. Although these systems implement and follow the FAIR principles to different extents, there is a lack of model and metadata interoperability, as well as a lack of integration with broader MLOps techniques.
Besides domain-specific systems, there are large-scale general-purpose model repositories, like OpenML [11] (an open platform to share datasets and AI/ML models), Hugging Face [92], TensorFlow Hub (now integrated in Kaggle) or Replicate [74]. Although these repositories are focused on sharing, model metadata is often unstructured and limited, and most model repositories are detached from the training process of the model, thus making users unable to assess the provenance or trustworthiness of a given asset.
2.1.3 Model deployment and MLOps
The deployment of AI/ML models into production services is a critical step to move ahead from research into real-world applications. However, the current landscape of MLOps tools and practices [9] is very fragmented, with most solutions designed for industry needs, rather than research.
Online AI/ML model serving can be implemented by leveraging a large set of lightweight solutions, allowing users to deploy a model behind an API. General-purpose serving can be directly implemented via Fast API [73], which is widely used as a framework to build REST APIs. However, these solutions are normally developed ad-hoc and lack interoperability. DEEPaaS API [37] allows users to offer a common interface for AI/ML models by exposing their Python code via an HTTP endpoint. TensorFlow Serving allows to serve TensorFlow models via gRPC and HTTP endpoints. BentoML allows users to package the models along with their dependencies into a deployable and self-contained asset. High-performance serving and inference can be implemented via more specific systems, like the NVIDIA Triton Inference Server (supporting multiple AI/ML frameworks and model formats, including TensorFlow, ONNX or PyTorch). More specifically, vLLM [53] has emerged as a specialized and optimized tool to serve large language models (LLMs).
MLOps platforms such as Kubeflow [52] or Seldon Core [83] provide an industry-grade and robust framework to automate model deployment and monitoring, including continuous integration and delivery (CI/CD). However, this strong focus on industry-driven workflows overlooks the unique needs of scientific research, including interoperability, reproducibility, provenance tracking, and FAIR compliance. For example, Kubeflow supports scalable deployments on Kubernetes, but it lacks interoperability with other cloud management frameworks or the possibility to easily execute a given workload on a different cluster technology (like in High Performance Computing systems).
2.1.4 Open science, FAIR principles
The FAIR principles have become one of the main pillars of open-science practices, with significant investments in methods, tools and best practices in order to broader their adoption. As a matter of fact, the FAIR principles underpin the evolution of the EOSC, which, as already stated, aims to facilitate the “Web of FAIR Data and Services for Science”, incentivizing also broader open science practices by improving the research assessment system [24].
@@ 2.1.4 Open science, FAIR principles
In the context of AI/ML research, these aspects should not be considered just as a mere set of best practices, but as an important vehicle towards AI/ML model transparency, reproducibility and trustworthiness. The FAIR principles, when applied to data leveraging platforms like Zenodo or Dataverse, can address data-related risk factors for AI (such as data provenance or licensing) [36]. However, these platforms are not able to respond to the dynamic nature of AI/ML model development and deployment, hence limiting the FAIR principles adoption beyond data. Researchers that want to ensure the FAIRness of their complete AI/ML solutions must bundle together a set of tools, standards and practices, at the cost of reproducibility, efficiency of their research, or compliance with open science principles.
@@ 2.2 AI4EOSC motivation
In this fragmented context, there is an opportunity to implement an open-source solution that can address the AI/ML lifecycle end-to-end, from development to deployment, following the open science and FAIR principles by design, therefore giving a more comprehensive overview of all the processes involved, thus fostering reproducibility. That platform should be approachable enough for non-technical scientists to use and offer a ready-to-use environment and workflows to streamline AI model creation. The resulting models should be reusable and not tied to the platform that created them. Besides, in the current ecosystem of European Research Infrastructures and the EOSC Federation, such a platform must be federated and dynamic in nature, allowing to integrate a wide variety of resources, from disparate providers.
The AI4EOSC platform addresses these challenges through its federated architecture, providing a seamless platform to help scientists to develop, train and deploy AI models in a user-friendly manner, following the open science and the FAIR principles by design. AI4EOSC operationalizes open science for AI/ML research, providing a unified, end-to-end platform, that implements open science principles to modern, data-driven research workflows.
The platform presented in this work, whose architecture will be described in Section 3, is an evolution of the DEEP platform [56]. This platform consisted on three different systems, all connected to a shared set of storage and data services: The DEEP training facility, that users could leverage to train their models; the DEEP Open Catalog, a marketplace where uses could share the models; and the DEEP as a service, allowing to deploy ML models on a serverless architecture and query them though a standardized model API.
Building on the basis of the former DEEP platform, AI4EOSC aims to solve the specific challenges that have been identified in this section (among others), through the whole AI/ML lifecycle:
@@
    Research-first and FAIR-by-design platform, prioritizing reproducibility, collaboration, and transparency, with rich metadata for datasets, models, and workflows, ensuring that all research outputs are findable, accessible, interoperable, and reusable without placing additional burdens on researchers.
@@
    MLOps practices tailored for research, extending MLOps principles to support reproducibility, provenance, and FAIR compliance by design via platform-level and user-defined CI/CD pipelines, integrated with data and model monitoring tools.
@@
    End-to-End provenance metadata for AI/ML models, capturing the full lineage of AI/ML processes, from data preprocessing to model deployment, including resource utilization. AI4EOSC leverages standardized provenance models (W3C PROV), ensuring that every step is traceable and reproducible.
@@
    Semantic (via MLDCAT-AP) and technical interoperability with other platforms, such as AI4Life, ensuring seamless interactions with other marketplaces, making it possible to implement and extend these integration towards other technologies.
@@
    Seamless EOSC Integration: AI4EOSC is natively integrated with the EOSC, enabling researchers to leverage EOSC’s infrastructure for data sharing (e.g., EOSC EU Node resources), collaboration, and long-term preservation. This integration ensures that AI/ML research is not only FAIR but also interoperable with the broader European research ecosystem.
@@
    Unified development environment with extensibility and flexible customization methods, allowing to easily integrate additional tools, like image annotation or federated learning.
@@
    Federated workload management system, being able to scale across many physically distributed datacenters and e-Infrastructures, with a variety of deployment and execution methods.

### 05 May 2026 v3

@@ Abstract
The rapid growth of Artificial Intelligence and Machine Learning in scientific research has highlighted a gap between industry-standard MLOps tools and platforms, and the unique requirements of modern and Open Science, particularly regarding the FAIR (Findable, Accessible, Interoperable, and Reusable) principles. This paper presents AI4EOSC, a federated, open-source platform designed to operationalize the full AI/ML lifecycle within the European Open Science Cloud (EOSC) ecosystem. Our methodology tackles the fragmentation of distributed research infrastructures by integrating a modular and distributed architecture comprising an AI development platform, a serverless AI-as-a-Service layer, and a federated orchestration model that is able to integrate heterogeneous compute and storage resources from distributed e-Infrastructures. AI4EOSC also introduces a “FAIR-by-design” approach that enforces metadata standardization (via MLDCAT-AP) and W3C PROV-compliant provenance tracking through a platform-integrated CI/CD pipeline. AI4EOSC added value is demonstrated through the delivery of a diverse set of community installations, showing consistent and seamless deployment across heterogeneous cloud providers. These installations are validated by a set of scientific cases, showing how our work reduces the manual burden on researchers while ensuring high levels of reproducibility and interoperability and providing an unified environment for development, training, and production of AI/ML models in the EOSC.
@@
Cloud Platform , Artificial Intelligence , Machine Learning , European Open Science Cloud , MLOps , Federated Learning , Provenance Tracking , W3C PROV , Open Science Infrastructure
@@ 1 Introduction
The adoption of Artificial Intelligence (AI) and Machine Learning (ML) in scientific research has moved beyond early experimentation. The impact of these techniques over the last years has revolutionized the field in a wide range of applications —e.g., by making it possible to tackle new scientific endeavors [19, 34, 25], by automatizing manual and tedious tasks [35] or by speeding up some time-consuming and effort-intensive tasks [55], just to cite some examples. These techniques are a reality, and they form nowadays part of the basic set of tools for a wide set of scientists and researchers [32]. However, this rapid adoption has exposed a critical gap: the infrastructure and practices needed to develop, share, and deploy AI/ML models in a reproducible, transparent and interoperable manner are still missing, or remain fragmented, within the scientific community.
This gap can be identified across two different axes. On the one hand, there is a large lifecycle fragmentation: the tools that are available for researchers working on AI/ML (integrated development environments, experiment tracking, model repositories, compute infrastructures, etc.) are normally independent components that work in isolation, each of them addressing a different AI/ML lifecycle phase. Assembling all these components together, so that they can support a researcher through the development of their pipelines, poses a significant technological burden on individual scientists. On the other hand, from the scientific standpoint, these tools have been developed without the specific needs of scientific research, and in particular there is a clear absence of FAIR-by-design enforcement. The FAIR principles (that stand for Findable, Accessible, Interoperable, Reusable) [58] have become a key pillar for open science practices, and nowadays are the norm in scientific research [^1] The FAIR principles aim to provide guiding principles and standards designed to optimize the reuse of digital assets. However, in spite of this fact, their application on AI/ML assets remain largely manual and hence its adoption is uneven.
We argue that these two axes cannot be treated as independent problems that can be tackled separately. General data platforms address data-level FAIRness, but they are not well suited to the dynamic nature of model development, where provenance, metadata, and deployment artifacts evolve continuously. FAIR compliance of AI/ML assets cannot be easily achieved by just adding metadata on top of a fragmented pipeline, and structural (i.e., FAIR-by-design) enforcement is needed. Metadata generation and their interoperability is only easily achievable if a single platform controls how assets are generated, built, versioned, and published; provenance records are only meaningful if they can capture the whole chain from data ingestion through training to deployment; interoperability cannot be achieved if there are vendor lock-ins or other trade-offs. This creates a significant architectural dependence, meaning that end-to-end integration is a necessary requirement in order to achieve a FAIR-by-design enforcement. To the best of our knowledge, there is not an open-source platform that addresses both axes together, particularly within the federated, multi-institutional context of European research infrastructures.
Machine Learning Operations (MLOps) is the closest analogue from the industry. MLOps has emerged as a set of practices and tools designed to ease the management of the entire AI/ML lifecycle [8], analogous to how DevOps bridges the gap between software development and IT operations. While industrial MLOps platforms have been successfully adopted they are are designed for cluster-locked and vendor-controlled environments, lacking native federation across heterogeneous and distributed e-Infrastructures. Moreover, its adoption in the scientific and research context is not that prominent and the systems supporting it prioritize operational performance over the scientific transparency, open standards, FAIR compliance that research demands.
@@
This gap is particularly critical in the context of the European Open Science Cloud (EOSC) ecosystem. The EOSC is an European Commission initiative aimed at creating the “web of FAIR data and services”, as a transformative process towards the promotion of open science practices among researchers, also delivering a comprehensive environment for the sharing, processing, and analysis of research data [21]. In the context of AI/ML applications and in the EOSC ecosystem, the heterogeneity of compute providers and the diversity of scientific communities makes more complex to reach this ambitious vision, and tackling it from a simple infrastructure standpoint is not enough, requiring of platforms that can operationalize open science principles across the entire AI/ML lifecycle.
In this context, we present the AI4EOSC platform, a federated open-source platform that tackles both axes of the gap simultaneously. The central architectural contribution of this work is the integration of the full AI/ML lifecycle (going from interactive development environments and federated training, through serverless deployment and drift monitoring, to W3C PROV-compliant provenance tracking and semantic interoperability via MLDCAT-AP) into a unified system where FAIR principles are enforced at the infrastructure level, automatically and transparently, without placing additional burden on researchers. AI4EOSC is tightly integrated within the EOSC ecosystem, spans the cloud-to-edge continuum, integrates natively with EOSC research infrastructures, and has been deployed and validated in production by three independent scientific communities.
@@
The remainder of this manuscript is structured as follows: Section 2 introduces the background and related work. In Section 3 and Section 4, we discuss the general architecture of the AI4EOSC platform and the design considerations that have underpinned its development. In Section 5 we provide performance metrics that validate the robustness of the platform. In Section 6 we explain how different scientists can leverage the platform across the full AI/ML lifecycle. Finally, Section 7 draws out the main conclusions and future work.
@@ 2 Related work and motivation
While there are many private and proprietary platforms to deploy AI/ML workloads, few have focused on addressing the specific open science and reproducibility requirements key to scientific research. The current landscape is characterized by framework fragmentation, where existing solutions address specific parts of the AI/ML lifecycle, often in isolation. We argue that there is a clear lack of integrated, open-science-focused solutions that span the entire lifecycle from development to deployment.
2.1 ML lifecycle platforms and Industrial MLOps
Standard industrial or enterprise frameworks like Kubeflow [36] or Polyaxon [46] provide robust automation but remain largely “cluster-locked” to specific Kubernetes environments. Federation of these clusters is not possible without the usage of additional tools or specialized control planes. This restricts their utility for executing workloads across the geographically distributed research e-Infrastructures typically used in scientific collaborations. Furthermore, while some the existing tools offer experiment tracking (either open source like MLflow [60] or LabML [33], or proprietary ones like Weights and Biases [22], Valohai [56] or Comet [20]), their focus remains on industrial workflows with limited support for open science and automated FAIR (Findable, Accessible, Interoperable, and Reusable) principles.
@@ 2.1 ML lifecycle platforms and Industrial MLOps
Moreover, end-to-end support for the whole AI/ML lifecycle process exists as commercial and closed source products (e.g., H2O.ai [31]) and restricts model deployment or the data storage to their specific cloud solutions (like Google Vertex AI [4], Amazon Sage Maker [5] or Azure ML Studio [39]), thus defeating the open principles on which science is built. Similarly, model serving solutions like Seldon Core [54] prioritize operational performance over the scientific transparency and deep tracking required for research trustworthiness.
2.2 Model Repositories and Scientific Catalogs
General-purpose repositories like Hugging Face [59], OpenML [10] (an open platform to share datasets and AI/ML models) and specialized scientific catalogs like the BioImage Model Zoo [41] or Kipoi [6] have demonstrated as essential for sharing assets with a broader community. However, model metadata is often unstructured and limited and the repositories are often detached from the training process, making it impossible for users to assess the full lineage or provenance of a model. AI4EOSC overcomes this by integrating the repository directly with a platform-enforced CI/CD provenance pipeline, ensuring users can fully assess the provenance of any given asset.
2.2.1 Provenance in ML Pipelines
The AI4EOSC provenance system shares its core objective with other PROV-based ML pipeline initiatives, most notably yProv [42], which also treats provenance as a first-class concern and aims to capture end-to-end relationships between data, models, and processing steps. However, the two approaches reflect a fundamental design trade-off. yProv achieves fine-grained lineage tracking by integrating tightly with the execution environment, instrumenting pipeline components to emit provenance information at runtime. This enables detailed capture of model evolution, parameter tuning, and intermediate data transformations, making it well suited to controlled settings where the full pipeline is under unified management.
AI4EOSC takes the opposite approach: provenance is collected non-intrusively from metadata already exposed by existing platform components, without requiring any modification to their internal behavior. While this constrains the level of detail to what those components natively provide, it offers significantly greater flexibility and ease of adoption in heterogeneous, real-world deployment environments where pipeline components are diverse and independently operated. Furthermore, the AI4EOSC approach explicitly targets semantic interoperability by producing PROV-O compliant JSON-LD output and supporting alignment with domain-specific ontologies, enabling provenance information to be reused and queried across system boundaries rather than remaining confined to the originating pipeline.
2.3 Evolution from the DEEP Platform
The DEEP-Hybrid-DataCloud (DEEP) [38] platform pioneered a distributed architecture for transparent access to European e-Infrastructures with focus on scientific applications. While DEEP successfully provided standardized APIs (DEEPaaS) and eased access to accelerators (GPUs), it lacked the deep semantic interoperability and automated metadata generation now required within the EOSC federation and ecosystem. AI4EOSC represents a significant evolution, introducing a serverless AI-as-a-Service (AIaaS) layer, federated learning (FL) capabilities, and a comprehensive FAIR-by-design framework that was previously absent.
2.4 Open Science and FAIR Principles
@@
The FAIR principles have become a central pillar of modern open-science practices. In the context of AI/ML research, these aspects should not be considered just as a mere set of best practices, but as an important vehicle towards AI/ML model transparency, reproducibility and trustworthiness. While general data platforms like Zenodo or Dataverse address data-related FAIRness (such as data provenance or licensing) [26], they are often unable to respond to the dynamic nature of AI/ML model development and deployment, hence limiting the FAIR principles adoption beyond data. Researchers that want to ensure the FAIRness of their complete AI/ML solutions must bundle together a set of tools, standards and practices, at the cost of reproducibility, efficiency of their research, or compliance with open science principles. AI4EOSC addresses this by operationalizing these principles across the entire AI/ML lifecycle through automated, “FAIR-by-design” mechanisms, that are easily extensible in order to accommodate new or unforeseen requirements.
@@ 2.5 AI4EOSC motivation
In this fragmented context, AI4EOSC addresses the need for an approachable, open-source solution that streamlines model creation for both technical and non-technical scientists. The platform differentiates itself through several core architectural innovations (as described in Section 3):
@@
    Transparent Inter-Cluster Federation: Our federated workload management system is able to scale across many physically distributed datacenters and e-Infrastructures, with a variety of deployment and execution methods, via a single and unified control plane that avoids the infrastructure silos of standard MLOps stacks.
@@
    Cloud-to-Edge Continuum: AI4EOSC mobilizes resources across a continuum, enabling deployment on heterogeneous hardware including low-capacity edge devices and cloud computing providers.
@@
    Hardened Federated Learning support: The platform extends standard Federated Learning (FL) with token-based authentication, novel aggregation methods, divergence monitoring and server-side differential privacy and metric privacy notions to defend against client inference attacks.
@@
    Integrated Provenance Tracking: The platform automatically transforms training fragments into W3C PROV-compliant RDF graphs, providing lineage traceability that is currently detached or proprietary in industry platforms.
@@
    Semantic Interoperability: AI4EOSC leverages MLDCAT-AP in order to deliver interoperability at the semantic level with other platforms and data spaces, enabling for seamless interactions with other marketplaces.
@@
    Integrated Drift Monitoring: Through the Frouros library and the DriftWatch system, researchers can detect and visualize concept and data drift at inference time.
@@
    Seamless EOSC Integration: AI4EOSC is natively integrated with the EOSC, enabling researchers to leverage EOSC’s infrastructure for data sharing (e.g., EOSC EU Node resources), collaboration, or long-term preservation. This integration ensures that AI/ML research is not only FAIR but also interoperable with the broader European research ecosystem.
To synthesize these differences and explicitly contrast AI4EOSC with the state-of-the-art, Table 1 provides a feature-wise comparison across key technical and scientific dimensions.
Kubeflow Polyaxon Seldon SageMaker OpenML Hugging DEEP AI4EOSC
Feature [36] [46] [54] [5] [10] Face [59] [38] (This work)
Scientific Focus Low Low Low Low High Medium High High
Openness Full Open-core Open-core Closed Full Partial Full Full
FAIR-by-Design No No No No Partial No Low Native
Provenance Limited Limited Limited Proprietary Limited Limited Low W3C PROV
LLM Integration Partial Partial Partial High No High No Native (RAG)
Learning Schemes Dist. Dist. N/A Dist. N/A N/A Dist. Fed. & Dist.
Drift Detection External Limited Yes Yes No No No Native
Infrastructure Kubernetes Kubernetes K8s/Cloud Prop. Cloud N/A N/A e-Infras e-Infras/Edge
Federation Cluster-lock Cluster-lock Cluster-lock Vendor-lock N/A N/A Partial Inter-cluster
Delivery Model Software Both Both Platform Both Platform Both Both
EOSC integration No No No No No No Partial Native
Table 1: Feature-wise comparison of AI4EOSC against representative industrial MLOps stacks, scientific tools, and the predecessor DEEP platform.